Find the molar mass of a juice drink that contains 1000 mg of Vitamin C in 757 g of water. The change in freezing point depression is 0.014 ⁰C.  Kf = 1.86 °C/m       

 


1
Expert's answer
2021-10-22T03:53:08-0400

The freezing point depression:

"∆T = k_f \times c_m"

kf = the molal freezing point depression constant

cm = the molality of the solute

"k_f = 1.86 \\

c_m = \frac{n}{m_{water}} = \frac{\frac{m}{M}}{m_{water}} \\

0.014 = 1.86 \times \frac{\frac{1 \;g}{M}}{0.757 \;kg} \\

0.014 = \frac{1.408}{M} \\

M = 100.57 \;g/mol"